A motorcycle was observed to have a velocity of 10m/s when it passed by town A,
and a velocity of 25m/s when it reached town B. If the two towns are 10km apart,
what was the motorcycle’s acceleration, as well as its travel time?
1
Expert's answer
2021-09-17T09:22:56-0400

The distance traveled by motorcycle 

"d=\frac{v_2^2-v_1^2}{2a}"

The acceleration

"a=\frac{v_2^2-v_1^2}{2d}"

"a=\rm\frac{(25\: m/s)^2-(10\: m/s)^2}{2*10000\: m}=0.026 \: m/s^2"